OpenStudy (anonymous):

Find the difference. k / k^2 - 9k + 20 - 5 / k^2 - 9k + 20 a. k - 5 / k - 4 b. 1 / k + 4 c. 1 / k - 4 d k - 5 / 0 Help me please.

OpenStudy (mathmale):

Your expression (that is, the function dependent upon the variable k) is ambiguous (it could be interpreted in different ways). Please make it crystal clear what you mean through liberal use of parentheses, so that no one needs to ask you for clarification. Without parentheses in this expression, the reader will waste time trying to determine whether you meant\[1) \frac{ k }{ k^2 }, ~2) \frac{ k }{ k^2-9k }, ~or~3) \frac{ k }{ k^2-9k+20 }\]

OpenStudy (anonymous):

the last one.

OpenStudy (mathmale):

Once you've used parentheses appropriately for greater clarity, or used Equation Editor (below) to present your expression, I'd be glad to help you with the problem itself.

OpenStudy (mathmale):

Re-write the WHOLE thing with parentheses, please.

OpenStudy (anonymous):

its cause the question and answers do not have parentheses..

OpenStudy (anonymous):

like that's how they're written, and i do not know how to write them in para..

OpenStudy (mathstudent55):

Do you mean this? \(\dfrac{k}{k^2 - 9k + 20} - \dfrac{5}{ k^2 - 9k + 20} \)

OpenStudy (anonymous):

exactly.

OpenStudy (mathstudent55):

Ok. You are subtracting fractions. What do you need to have to be able to add or subtract fractions?

OpenStudy (anonymous):

add the Ks,

OpenStudy (mathmale):

Soccergirl: You're missing my point completely. it is your responsibility to present problems in such a way that there is no ambiguity.

OpenStudy (anonymous):

okay mathmale. i've written problems like this like ten already no1 had a problem only you tbh,

OpenStudy (anonymous):

sorry mathmale . :c

OpenStudy (mathmale):

Have it your way. i tried to help you. If you don't want my advice, so be it.

OpenStudy (mathstudent55):

For example, can you subtract these two fractions the way they are written, or do you need to do something first? \(\dfrac{3}{4} - \dfrac{5}{7} \)

OpenStudy (anonymous):

must change the sign?

OpenStudy (anonymous):

is it B?

OpenStudy (mathstudent55):

No. If you need to subtract, you need the keep the minus sign. Do you see the denominators? They are different. What do you need to do to the different denominators to be able to subtract fractions?

OpenStudy (anonymous):

crap idk :C

OpenStudy (mathstudent55):

If you can't subtract simple fractions like \(\dfrac{3}{4} - \dfrac{5}{7}\), how can you expect to be able to subtract the fractions you were given that involve factoring which is more complicated than simple numbers?

OpenStudy (mathmale):

SoccerGirl: I'm continuing to monitor (watch) this conversation. It surely appears that you're trying to guess your way through this problem. In the long run, you'd be better off by far actually learning the material.

OpenStudy (anonymous):

I am not trying to guess. Stop trying to say that, I am actually trying to learn.

OpenStudy (anonymous):

Go through all my questions, do not say such a thing if you are not here to help leave my question please.

OpenStudy (mathmale):

Just a warning: SoccerGirl: You are spinning your wheels and getting nowhere. Please think this through and see whether you're willing and able to change your attitude.

OpenStudy (anonymous):

You're just mad cause I did not do it your way.

OpenStudy (anonymous):

Change your own attitude. Seriously leave my question.

OpenStudy (anonymous):

You're just arguing and not helping, leave before I report you for disorderly conduct.

OpenStudy (mathstudent55):

I'll tell you this much. To add or subtract fractions, you need a common denominator. The problem you were given already has a common denominator for the fractions, so just subtract the numerators and set it over the same denominator. Then factor the denominator, and see what you can do to reduce the fraction.

OpenStudy (mathmale):

More sad than mad, seeing that you don't appreciate advice from someone trying to help you. I have 43 years of experience in teaching math, so I do know a thing or two. Good bye and good luck, Soccer Girl: You need it.

OpenStudy (anonymous):

You're a teacher that kids bullet? Okay. Just leave my question.

OpenStudy (anonymous):

Okay my brother helped me, he said it was A. Is he wrong? If so please explain why,.

OpenStudy (mathmale):

Sorry, SoccerGirl: For tonight at least, I am out of this question (at your request).

OpenStudy (anonymous):

Alright Mathmale.

OpenStudy (mathstudent55):

Start with this: \(\dfrac{k}{k^2 - 9k + 20} - \dfrac{5}{ k^2 - 9k + 20}\) Since you have a common denominator, just subtract the numerators and set it over the common denominator: \(=\dfrac{k-5}{k^2 - 9k + 20} \) Do you follow so far?

OpenStudy (anonymous):

I do.

OpenStudy (anonymous):

So what I said was correct! :d

OpenStudy (mathstudent55):

Ok, now you need to try to reduce the fraction. To do that, you need to factor the denominator.

OpenStudy (anonymous):

I know i know, but i was right im proud :D

OpenStudy (mathstudent55):

\(=\dfrac{k-5}{(k - 5)(k - 4)}\) Now you divide the numerator and denominator by the same common factor (which people usually call cancelling). \(=\dfrac{\cancel{k-5}~~1}{\cancel{(k - 5)}(k - 4)}\) The final answer is: \(=\dfrac{1}{k - 4}\)

OpenStudy (anonymous):

AWH wtf i was wrong :C

OpenStudy (anonymous):

i thought i was right :( im sad now :/

OpenStudy (mathstudent55):

That means the answer is C.

OpenStudy (anonymous):

Thank you so much for your help! :) <3

OpenStudy (mathmale):

My warm thanks to you, @mathstudent55.

OpenStudy (mathstudent55):

You need to cover the material better, so you know how to solve the problem. Then you won't have to guess. Go over this problem again. There are several concepts here that you need to learn. 1. To add or subtract fractions, you need a common denominator. 2. You need to learn how to factor polynomials.

OpenStudy (mathstudent55):

@mathmale Thanks.

OpenStudy (anonymous):

@mathmale i understand your confusion but if the asker (in this case @Soccergirl52 ) does not know what it means herself then it would be easier to present it as is and try to figure out the best way to solve it even if it means you have to try again if you make a mistake.